OpenStudy (anonymous):

i just want to know the answer to the square root of 7 minus 3 over 4....

OpenStudy (anonymous):

\[\frac{\sqrt{7-3}}{4}=\frac{\sqrt{4}}{4}=\frac{2}{4}=\frac{1}{2}\]

OpenStudy (anonymous):

i'm sorry. the minus three is not under the square root sign. i don't know how to type a square root sign here.